OpenStudy (anonymous):

Solve the system of equations: what is the x-value for the solution? 2x + 7y = 3 and x = 1 - 4y

hero (hero):

2(1-4y) + 7y = 3 2 - 8y + 7y= 3 -y = 1 y = -1 x = 5